# HD 44892: The youngest (or oldest?) gas-harbouring debris disc around an intermediate mass star

### v1

@@ 4.2 UVES results
There is, however, some variability in the residual flux present in both Ca ii K and H lines (Figure 7), which is not present in the sodium lines. This variability does not match with any of the clouds in the line of sight nor the nearby comparison stars. While not at the radial velocity of the parent star, this variability in the residual absorption might be caused by secondary gas released from volatile bodies like in case of Pic (Dent et al., 2014) or a gas wind-disc, like the one found around NO Lup (Lovell et al., 2021b), although this cannot be confirmed with the data we currently have. Thus, it is possible we have detected variable, circumstellar gas around HD 44892.
Following the methods described in Fairlamb et al. (2017) we calculated the mass accretion rate of HD 44892 from the H spectrum. We found accretion rates of = -7.64 and = -7.69 for the two observed epochs, which are consistent with those of other Herbig Ae/Be stars (Wichittanakom et al., 2020). This further suggests that HD 44892 is a pre-MS star, with an estimated age of 2.1 Myr, as accretion is characteristic of young stellar objects.
@@ 5.1.1 Continuum
The continuum emission we have detected around HD 44892 is only marginally resolved along the major axis. Because of this, we are unable to estimate the orbital distance to the debris disc using our current data, and hence cannot determine the dust temperature, on which dust mass depends. To provide a more reliable estimate of the dust mass, we have utilised the ‘simalma’ task, which is part of CASA package. We first created model images of elliptical dust rings located at a range of orbital distances from the star (with disc radii extending to 20-100 au along the major axis), which we then input into ‘simalma’. The simulations were carried out using the exact same observational setup as our real continuum observations - including observing time, precipitable water vapour (pwv), frequency, antenna setup, and other parameters - in order to accurately replicate the noise, resolution, and observed flux.
We used the CASA task ‘imfit’ to compare the simulated images to the actual data, and found that a ring located approximately 40-45 au from the star best reproduced the observed emission. Since the minor axis is not resolved, we cannot give an estimate to its size. Based on this distance, we estimated the dust temperature using Equation 2:
@@
where is the source emission, is the distance to the source, is the Planck function (dependent on dust temperature, ), and is the dust opacity, which is very uncertain. Here, we assume (Beckwith et al., 1990), where is the frequency of observations. This results in a minimum dust mass at the dust temperature of 115 K of 0.019 0.009 M⊕. This mass is highly dependent on the dust temperature, which is uncertain given the emission is only marginally resolved along the major axis. For comparison, a lower dust temperature of 50 K would yield a higher dust mass estimate of 0.05 0.03.
Based on the expected lifetime of the observed dust, we can estimate whether it is primordial (left over from system formation) or secondary (released in planetesimal collisions). In the absence of gas, the lifetime of millimetre dust would be dominated by grain-grain collisions. The longest timescale at which this amount of millimetre dust can be destroyed, assuming collisions with dust of the same size, is if we distribute it in the entire region up to 40-45 au. In this scenario, following Pearce et al. (2021) (Equation 8, based on Wyatt & Dent, 2002) we obtain the dust lifetime of order Myr, which is much smaller than the estimated age of HD 44892. Whilst the derived lifetime could be extended in the presence of gas, which can damp grains and reduce their impact velocities, very significant damping would be required to extend the lifetime up the the star age. Hence the estimated lifetime, combined with the relatively low dust mass compared to protoplanetary discs (Figure 2), leads us to believe that at least some of the dust is secondary in nature. This is typical for hybrid discs, which are thought to comprise primordial gas and second-generation dust (e.g. Kóspál et al. 2013; Moór et al. 2013).
@@ 5.1.2 Gas
To calculate the CO gas mass in this disc we follow methods described in Miley et al. (2019) and Matrà et al. (2015). Since our detection is in 12CO (2-1) only, we use Equation 4 to calculate the minimum CO gas mass, assuming local thermodynamic equilibrium:
@@
Given the young age of HD 44892 and its transitional stage from a protoplanetary to a debris disc, the primordial origin of gas cannot be ruled out (Nakatani et al., 2021). Primordial gas is leftover from the protoplanetary stage, where it is dominated by the ISM-like abundances of hydrogen (H) and carbon monoxide (CO), with other gases contributing to the total gas mass. Assuming the disc is hybrid, we use the ISM ratio of /H2= to estimate the total gas mass in this disc, under the assumption of a substantial hydrogen presence in the disc. If the gas is indeed primordial, we estimate the H2 mass, and hence total disc gas mass, to be 0.06 0.02 for 20 K gas and 0.12 0.01 for 50 K gas.
From the moment 0 map made for the two non-detections, 13CO and C18O, we derived upper limits to the CO gas mass. By overlaying the 3 contours of 12CO emission on this map, we derived the RMS within this region to be 30.28 mJy km s-1, with 3 mJy km s-1, which we use as 13CO and C18O flux for the purpose of gas mass calculations, since the dataset is common between them. We adopt the ISM-like relative ratios of CO/13CO and CO/C18O as 77 and 550, respectively (Wilson & Rood, 1994). Using Equation 4 with adjusted values for and , we obtained the upper limits to the CO gas mass as presented in Table 1. However, in debris discs there typically is more 13CO than C18O (see, e.g., Miley et al., 2018), hence 13CO gives a more reliable estimate of CO gas mass upper limit. The total CO gas mass in the debris disc of HD 44892 should hence be contained between the lower limits from the 12CO emission and upper limits from 13CO non-detection.
@@ 5.1.3 Gas and dust emission morphology
Examining the structure of both the continuum and 12CO emission (Figure 8), we can see that the dust emission clearly has a different morphology than the 12CO emission. This direct comparison in their morphology is possible because of the similar signal-to-noise ratio of both dust and 12CO emission. The 12CO integrated emission is not resolved, therefore it follows the beam distribution. As mentioned in Section 3.2.1, the continuum emission is marginally resolved, which can be further seen in Figure 9. Here, the longest UV-distance baselines, at 800 m, deviate from a flat profile, which indicates the continuum emission is not unresolved, since a point-like source would have a constant amplitude across all baselines. Since the extended continuum structures are resolved, we can see that the dust emission is asymmetrical and has a more complex structure than CO emission - it seems to extend more to the north-west of the host star. The larger extent of continuum emission is atypical in protoplanetary discs, and more common in debris discs (e.g., Inutsuka et al., 2023; Moór et al., 2019; Di Folco et al., 2020).
@@ 5.2 SED
Figure 10 shows the SED of the HD 44892 debris disc, assuming optically thin dust, which is best modelled with a single-temperature belt structure. Due to the uncertainties in the temperature of the parent star Teff = 7500 500K, we present the SED fit to the stellar temperature of 7500 K along with the alternative minimum and maximum values of 7000 K and 8000 K. The fit to the excess is highly dependent on the assumed stellar temperature, changing the resultant dust-belt temperatures. The best-fit SED (left plot of Figure 10, 7000 K) suggests that the debris disc is composed of a warm belt at 739 K, based on which we estimate the location of this component belt using Equation 2.
We estimate the distance of the 739 K dust belt to be 1.1 au from the parent star. The two higher-temperature fits produce additional excess points that cannot be fitted with the existing blackbody component and would hence require an additional, even hotter belt fit. However, at such high temperatures, around A-type stars, it would be physically impossible for the dust to survive, suggesting the stellar fit might not not be accurate. Carbon grains 0.1 au away from A-type stars sublimate at 2000 K (e.g., Lebreton et al., 2013; Pearce et al., 2020). Silicates sublimate at comparatively cooler temperatures of 1200 K (e.g., Sezestre et al., 2019).
In addition to the SED data shown in Figure 10, there exists a 100m upper limit listed in the IRAS catalogue. However, upon examination, this emission seems to originate from an extended background source unrelated to HD 44892, hence we did not include it in the SED fit.
@@ 5.3 IRS
Using data from the IRS spectrograph (the Infrared Spectrograph on the Spitzer Space Telescope), in Figure 11 we examined and compared the mid-IR spectrum of HD 44892 (PI Bouwman, ID 3470) to that of protoplanetary discs HD 97048 (PI Houck, ID 2) and HD 36112 (PI Bouwman, ID 3470), and debris discs with gas HD 138813 (PI Carpenter, ID 30091) and HD 131835 (PI Houck, ID 40651). The spectrum of HD 44892 is quite complex and looks very different to those of both protoplanetary and debris discs, and it shows a distinctive feature at the wavelength of 12 m, which is not present in the spectra of other objects. The comparison shows that HD 44892 is unique and likely caught at the transition from protoplanetary to debris disc stage.
@@ 6 HD 44892 in the context of known gas-harbouring debris discs
In Figure 12 we compare the dust and CO gas masses we calculated for HD 44892 with Herbig star discs (flux adapted from Moór et al., 2017) and all other debris discs with CO gas mass measurements to date (see Table 2 for detailed list), highlighting the discs HD 141569 (Miley et al., 2018), NO Lup (Lovell et al., 2021a), Fomalhaut (Matrà et al., 2017b; MacGregor et al., 2017), and Crv (Marino et al., 2017). HD 44892, due to its probable very young age, populates an area where we would expect protoplanetary discs to reside, but both its dust and CO gas masses are much lower than expected for PPDs.
@@
Just like in the case of HD 141569, the gas in HD 44892 exhibits an apparently continuous morphology, albeit severely affected by the fact that the 12CO emission is spatially unresolved. On the other hand, even with limited spatial resolution, we see asymmetry in dust distribution, a very prominent feature of HD 141569 (Biller et al., 2015), and dust is present out to a radius beyond the 12CO gas detection, like in HD 141569 (Miley et al., 2018).
At the young age of HD 44892, apart from HD 141569, only one other comparable 12CO detection in a debris disc exists - that in the disc around NO Lup, a 0.7 star with an age estimate of 1-3 Myr. The two other possible 12CO detections presented in Lovell et al. (2021a), around Sz 108B and Lup 818s, are likely caused by CO cloud (ISM) gas instead. For NO Lup, using continuum and 12CO fluxes from Lovell et al. (2021a), we derive the CO gas mass, assuming gas temperature of 50 K, of (5.03 1.13) 10-5 . At the reported dust temperature of 20 K, we estimate the dust mass in the disc to be (4.00 1.38) 10-2 . These values are comparable to those we derive for HD 44892. HD 44892, similarly to NO Lup, has lost its protoplanetary disc, but retains some gas and dust. However, given their very different stellar masses and structure, the mechanisms responsible for dispersal are very likely different, i.e., FUV/EUV photoevaporation for the radiative A-type star HD 44892, and X-ray photoevaporation for the convective low mass star NO Lup (Kunitomo et al., 2021).
@@
Only three debris discs with gas have been found around stars older than 100 Myr: HD 23642 (V* V1229 Tau) at 170 Myr (Péricaud et al., 2017; Southworth et al., 2023), Fomalhaut (HD 126956) at 440 Myr (Matrà et al., 2017b), and the 1-2 Gyr Crv (Marino et al., 2017). The older age possibility of HD 44892, at 800 Myr, would make it comparable to the latter two of these objects. For Fomalhaut, using data from Matrà et al. (2017b) and MacGregor et al. (2017), we derive a 50 K dust mass of and a CO gas mass of . For Crv we derive a dust mass at 50 K of and a CO gas mass of (with flux data from Marino et al., 2017). Both systems contain at least three orders of magnitude less CO gas mass than HD 44892 (see Figure 12). This further shows that if HD 44892 was 800 Myr, it would be an outlier, containing significantly high amounts of CO gas.
It is worth noting that discs exist around post-asymptotic giant branch (AGB) binary star systems, with dust masses even up to 1 (Kluska et al., 2022). These, however, are created from the mass-loss of primary object during the last stages of its evolution. HD 44892, at the age of 2.1 Myr, is approaching the zero-age main sequence (ZAMS) or, at the alternative age of 800 Myr, is crossing the Hertzsprung-Russell (HR) gap after leaving the MS. Both its stellar temperature and luminosity are inconsistent with the star being on the AGB. Moreover, there is no evidence to suggest that HD 44892 is a binary system, hence this scenario is inapplicable here.
The mass of debris discs decreases over time due to the collisional evolution, as the smallest grains are constantly removed from the system by radiation pressure from the star (Bonsor & Wyatt, 2010). This process, also called collisional cascade, causes larger bodies to collide and break down into smaller bodies, eventually creating dust. The typical lifetime of dust in debris discs is of the order 1 Myr (Krivov et al., 2006), meaning that this dust must be continuously replenished through subsequent collisions of planetesimals to remain detectable (Wyatt, 2008). Only the most massive debris discs are hence expected to remain detectable at advanced ages, and their excess emission is relatively low, as seen from Figure 2.

### v2

@@ 4.2 UVES results
There is, however, some variability in the residual flux present in both Ca ii K and H lines (Figure 8), which is not present in the sodium lines. This variability does not match with any of the clouds in the line of sight nor the nearby comparison stars. While not at the radial velocity of the parent star, this variability in the residual absorption might be caused by secondary gas released from volatile bodies like in case of Pic (Dent2014MolecularDisk) or a gas wind-disc, like the one found around NO Lup (Lovell2021RapidDisc), although this cannot be confirmed with the data we currently have. Thus, it is possible we have detected variable, circumstellar gas around HD 44892.
Following the methods described in Fairlamb2017ALines we calculated the mass accretion rate of HD 44892 from the H spectrum. We found accretion rates of = -7.64 and = -7.69 for the two observed epochs, which are consistent with those of other Herbig Ae/Be stars (Wichittanakom2020TheStars). This further suggests that HD 44892 is a pre-MS star, with an estimated age of 2.1 Myr, as accretion is characteristic of young stellar objects. Variability is also observed in the H line (Szewczyk et al., in prep), similar to that seen in the Ca ii K & H lines. Hence, the derived accretion rate may correspond to a temporary phase of high accretion and should therefore be regarded as an approximate estimate.
@@ 5.1.1 Continuum
The continuum emission we have detected around HD 44892 is only marginally resolved along the major axis. Because of this, we are unable to estimate the orbital distance to the debris disc using our current data, and hence cannot determine the dust temperature, on which dust mass depends. To provide a more reliable estimate of the dust mass, we have utilised the ‘simalma’ task, which is part of CASA package. We first created models, each containing a circular dust ring at various disc inclinations, located at a range of orbital distances from the star (with disc radii extending to 20-100 au along the major axis), which we then input into ‘simalma’. Because the continuum emission is only marginally resolved, we did not consider the ring width as a free parameter and fixed it to 1 px (0.2 au), small enough to not influence the fit. The aim of our approach was to determine the dust ring size that would best mimic our observations - effectively, the maximum radius at which the dust can be placed without the belt becoming resolved. The simulations were carried out using the exact same observational setup as our real continuum observations - including observing time, precipitable water vapour (pwv), frequency, antenna setup, and other parameters - in order to accurately replicate the noise, resolution, and observed flux.
We used the CASA task ‘imfit’ to compare the simulated images to the actual data, and found that a ring located approximately 40-45 au from the star best reproduced the observed emission. Since the minor axis is not resolved, we cannot give an estimate to its size. Based on this distance, assuming the dust particles behave as black bodies, we estimated the dust temperature using Equation 2:
@@
where is the source emission, is the distance to the source, is the Planck function (dependent on dust temperature, ), and is the dust opacity, which is very uncertain. Here, we assume [cm2 g-1] (Beckwith1990), where is the frequency of observations [Hz]. This results in a minimum dust mass at the dust temperature of 115 K of 0.019 0.009 M⊕. This mass is highly dependent on the dust temperature, which is uncertain given the emission is only marginally resolved along the major axis. For comparison, a lower dust temperature of 50 K would yield a higher dust mass estimate of 0.05 0.03.
Based on the expected lifetime of the observed dust, we can estimate whether it is primordial (left over from system formation) or secondary (released in planetesimal collisions). In the absence of gas, the lifetime of millimetre dust would be dominated by grain-grain collisions. The longest timescale at which this amount of millimetre dust can be destroyed, assuming collisions with dust of the same size, is if we distribute it in the entire region up to 40-45 au. In this scenario, following Pearce2021Fomalhaut (Equation 8, based on Wyatt2002Collisions) we obtain the dust lifetime of order Myr, which is much smaller than the estimated age of HD 44892. Whilst the derived lifetime could be extended in the presence of gas, which can damp grains and reduce their impact velocities, very significant damping would be required to extend the lifetime up the the star age. Hence the estimated lifetime, combined with the relatively low dust mass compared to protoplanetary discs (Figure 2), leads us to believe that at least some of the dust is secondary in nature. This is typical for hybrid discs, which are thought to comprise primordial gas and second-generation dust (e.g. Kospal2013ALMA21997; Moor2013ALMA21997).
@@ 5.1.2 Gas
To calculate the CO gas mass in this disc we follow methods described in Miley2019Asymmetric100546 and Matra2015COALMA. Since our detection is in 12CO (2-1) only, we use Equation 4 to calculate the minimum CO gas mass, assuming local thermodynamic equilibrium:
@@
Given the young age of HD 44892 and its transitional stage from a protoplanetary to a debris disc, the primordial origin of gas cannot be ruled out (Nakatani2021PhotoevaporationRemnants). Primordial gas is leftover from the protoplanetary stage, where it is dominated by the ISM-like abundances of molecular hydrogen (H2), with carbon monoxide (CO) along other gases contributing to the total gas mass. Assuming the disc is hybrid, we use the ISM ratio of /H2= to estimate the total gas mass in this disc, under the assumption of a substantial hydrogen presence in the disc. If the gas is indeed primordial, we estimate the H2 mass, and hence total disc gas mass, to be 0.06 0.02 for 20 K gas and 0.12 0.01 for 50 K gas. These estimates are based on the lower limit of CO gas mass (estimated from emission), hence the total gas mass in this disc may be higher.
From the moment 0 map made for the two non-detections, 13CO and C18O, we derived upper limits to the CO gas mass. By overlaying the 3 contours of 12CO emission on this map, we derived the RMS within this region to be 30.28 mJy km s-1, with 3 mJy km s-1, which we use as 13CO and C18O flux for the purpose of gas mass calculations, since the dataset is common between them. We adopt the ISM-like relative ratios of CO/13CO and CO/C18O as 77 and 550, respectively (Wilson1994AbundancesMedium). Using Equation 4 with adjusted values for and , we obtained the upper limits to the CO gas mass as presented in Table 1. However, since in debris discs 13CO is typically more abundant than C18O (see, e.g., Miley2018Unlocking141569), the upper limit on the CO gas mass derived from 13CO is more stringent. The total CO gas mass in the debris disc of HD 44892 should hence be contained between the lower limits from the 12CO emission and upper limits from 13CO non-detection.
@@ 5.1.3 Gas and dust emission morphology
Examining the structure of both the continuum and 12CO emission (Figure 9), we can see that the dust emission clearly has a different morphology than the 12CO emission. This direct comparison in their morphology is possible because of the similar signal-to-noise ratio of both dust and 12CO emission. The 12CO integrated emission is not spatially resolved, therefore it follows the beam distribution. As mentioned in Section 3.2.1, the continuum emission is marginally resolved, which can be further seen in Figure 4. Here, the longest UV-distance baselines, at 800 m, deviate from a flat profile, which indicates the continuum emission is not unresolved, since a point-like source would have a constant amplitude across all baselines. The extended continuum emission is only marginally resolved with an apparent asymmetry at a level of 4.5 , which may or may not indicate asymmetry in the dust disc. The larger extent of continuum emission with respect to CO emission is atypical in protoplanetary discs, and more common in debris discs (e.g., PPVII2023; Moor2019New32297; DiFolco2020). Higher resolution observations are necessary to ascertain the distribution of dust and CO gas in the HD 44892 system.
@@ 5.2 SED
Figure 10 shows the SED of the HD 44892 debris disc, assuming optically thin dust. The excess emission is best modelled with a single-temperature belt structure, fitted using a blackbody model. Due to the uncertainties in the temperature of the parent star Teff = 7500 500 K, we present the SED fit to the stellar temperature of 7500 K along with the alternative minimum and maximum values of 7000 K and 8000 K. The fit to the excess is highly dependent on the assumed stellar temperature, changing the resultant dust-belt temperatures. The best-fit SED (left plot of Figure 10, 7000 K) suggests that the debris disc is composed of a warm belt at 739 K, based on which we estimate the location of this component belt using Equation 2.
We estimate the distance of the 739 K dust belt to be 1.1 au from the parent star. The two higher-temperature fits produce additional excess points that cannot be fitted with the existing blackbody component and would hence require an additional, even hotter belt fit. However, at such high temperatures, around A-type stars, it would be physically impossible for the dust to survive, suggesting the stellar fit might not be accurate. Carbon grains 0.1 au away from A-type stars sublimate at 2000 K (e.g., Lebreton2013; Pearce2020). Silicates sublimate at comparatively cooler temperatures of 1200 K (e.g., Sezestre2019). In addition to the SED data shown in Figure 10, there exists a 100m upper limit listed in the IRAS catalogue, which we did not include in the SED fit, as it is very large and does not represent a meaningful constraint.
@@ 5.3 IRS
Using data from the IRS spectrograph444With reduced spectra obtained from https://cassis.sirtf.com (the Infrared Spectrograph on the Spitzer Space Telescope), in Figure 11 we examined and compared the mid-IR spectrum of HD 44892 (PI Bouwman, ID 3470) to that of protoplanetary discs HD 97048 (PI Houck, ID 2) and HD 36112 (PI Bouwman, ID 3470), and debris discs with gas HD 138813 (PI Carpenter, ID 30091) and HD 131835 (PI Houck, ID 40651). The spectrum of HD 44892 is quite complex and looks very different to those of both protoplanetary and debris discs, and it shows a distinctive feature at the wavelength of 12 m, which is not present in the spectra of other objects. The comparison shows that HD 44892 is unique and likely caught at the transition from protoplanetary to debris disc stage.
@@ 6 HD 44892 in the context of known gas-harbouring debris discs
In Figure 12 we compare the dust and CO gas masses we calculated for HD 44892 with Herbig star discs (flux adapted from Moor2017MolecularStars) and all other debris discs with CO gas mass measurements to date (see Table 2 for detailed list), highlighting the discs HD 141569 (Miley2018Unlocking141569), NO Lup (Lovell2021ALMADispersal), Fomalhaut (Matra2017DetectionComets; MacGregor2017ADisk), and Crv (Marino2017ALMAPlanets). HD 44892, due to its probable very young age, populates an area where we would expect protoplanetary discs to reside, but both its dust and CO gas masses are much lower than expected for PPDs.
@@
The accretion rate we derived for HD 44892 (see Section 4.2) is very similar to that measured for HD 141569 (Fairlamb2015) and is orders of magnitude higher than expected for CO-rich debris discs (10-11-10-13 , Kral2017PredictionsPlanetesimals; Borthakur2025). If we assume a secondary origin for the gas, where the gas is H-poor with no significant H2 reservoir, and adopting the upper limit on the total CO gas mass (based on the 13CO emission) of 10-2 , the CO in the disc would be depleted within just a few years. It is highly unlikely that a mechanism capable of replenishing such a large amount of secondary gas exists, thus the high accretion rate suggests that at least the inner disc is of primordial origin and dominated by H2 gas. However, given the accretion variability discussed in Section 4.2, this value may represent a phase of enhanced accretion. A more detailed discussion will be presented in Szewczyk et al. (in prep).
At the young age of HD 44892, apart from HD 141569, only one other comparable 12CO detection in a debris disc exists - that in the disc around NO Lup, a 0.7 star with an age estimate of 1-3 Myr. The two other possible 12CO detections presented in Lovell2021ALMADispersal, around Sz 108B and Lup 818s, are likely caused by CO cloud (ISM) gas instead. For NO Lup, using continuum and 12CO fluxes from Lovell2021ALMADispersal, we derive the CO gas mass, assuming gas temperature of 50 K, of (5.03 1.13) 10-5 . At the reported dust temperature of 20 K, we estimate the dust mass in the disc to be (4.00 1.38) 10-2 . These values are comparable to those we derive for HD 44892. HD 44892, similarly to NO Lup, has lost its protoplanetary disc, but retains some gas and dust. However, given their very different stellar masses and structure, the mechanisms responsible for dispersal are very likely different, i.e., FUV/EUV photoevaporation for the radiative A-type star HD 44892, and X-ray photoevaporation for the convective low mass star NO Lup (Kunitomo2021PhotoevaporativeStars).
@@
Only three debris discs with gas have been found around stars older than 100 Myr: HD 23642 (V* V1229 Tau) at 170 Myr (Pericaud2017TheDisks; Southworth2022ScutiHD23642), Fomalhaut (HD 216956) at 440 Myr (Matra2017DetectionComets), and the 1-2 Gyr Crv (Marino2017ALMAPlanets). The older age possibility of HD 44892, at 800 Myr, would make it comparable to the latter two of these objects. For Fomalhaut, using data from Matra2017DetectionComets and MacGregor2017ADisk, we derive a 50 K dust mass of and a CO gas mass of . For Crv we derive a dust mass at 50 K of and a CO gas mass of (with flux data from Marino2017ALMAPlanets). Both systems contain at least three orders of magnitude less CO gas mass than HD 44892 (see Figure 12). This further shows that if HD 44892 was 800 Myr, it would be an outlier, containing significantly high amounts of CO gas.
It is worth noting that discs exist around post-asymptotic giant branch (AGB) binary star systems, with dust masses even up to 1 (Kluska2022APlanets). These, however, are created from the mass-loss of primary object during the last stages of its evolution. HD 44892, at the age of 2.1 Myr, is approaching the zero-age main sequence (ZAMS) or, at the alternative age of 800 Myr, is crossing the Hertzsprung-Russell (HR) gap after leaving the MS. Both its stellar temperature and luminosity are inconsistent with the star being on the AGB. Moreover, there is no evidence to suggest that HD 44892 is a binary system, hence this scenario is inapplicable here.